### v1

@@ 3 Methods
The GLAT was developed following the established test development procedures outlined in Psychological and Educational Measurement (Thorndike et al., 1991). The development process involved: 1) creating a blueprint of relevant GenAI concepts, 2) generating an initial set of test items based on this blueprint, and 3) evaluating face and content validity through expert reviews and pilot studies, respectively. An item analysis was conducted using CTT, which involved selecting items based on item difficulty and discrimination index. The structural validity and reliability of the GLAT (RQ1) were assessed using IRT. The incremental validity of the GLAT (RQ2) was evaluated by analysing its effectiveness in predicting learners’ performance on tasks involving interaction with a GenAI chatbot, compared to a self-reported GenAI literacy instrument (Lee and Park, 2024). Details were elaborated in the following sections.
@@ 3.1 Participants
Three samples of higher education students were involved in various validation processes for the GLAT, following established standards (Thorndike et al., 1991; American Educational Research Association et al., 2014). The first validation study involved assessing the content validity and selecting the GLAT item using CTT. Responses were gathered from 200 higher education students (101 females). The second validation study aimed to evaluate the structural validity and reliability of the GLAT using IRT (RQ1). This study analysed responses from 355 higher education students (184 females). The final validation study focused on assessing the incremental validity of the GLAT (RQ2) and included 83 higher education students (46 females). All participants were recruited through Prolific [^1] https://www.prolific.co/, a reputable online research recruitment platform. Participants in the first and second studies received £1.5 for their time, while those in the final study were compensated £8 due to its increased complexity, as elaborated in Section 3.4. All studies were conducted using Qualtrics, with the item and option orders in the GLAT randomised (American Educational Research Association et al., 2014). Ethics approval was obtained from [Anonymised] University (Project ID: Anonymised), and informed consent was obtained from all participants.
@@ 3.2.1 Blueprint construction
The blueprint of the GLAT was developed based on the four dimensions of AI literacy proposed by Ng et al. (2021c), including 1) Know & Understand, 2) Use & Apply, 3) Evaluate & Create, and 4) Ethics, and focusing on the specific context of GenAI. To identify a set of relevant GenAI concepts in each dimension, we cover a wide range of resources, including academic publications in prestige journals (e.g., Nature and Science), reports and articles published by reputable organisations (e.g., UNESCO, MIT News, and Standford HAI), and education information released by leading GenAI technology companies (e.g., OpenAI, Google, Meta, and NVIDIA). The decision to include diverse sources, beyond traditional academic publications, was driven by the rapidly evolving landscape of GenAI. This encompasses foundational models (e.g., large language models and diffusion models) as well as supporting infrastructure and techniques (e.g., embedding databases and retrieval methods). Our process for extracting relevant concepts involved three steps. Initially, two researchers independently reviewed the source documents (n=19; links are available in the repository) and recorded pertinent GenAI concepts. Subsequently, they collaborated to consolidate similar concepts, resulting in a refined set of 25 concepts (see Table 1). A validation panel of three GenAI researchers then reviewed these concepts to ensure they provided reasonable coverage of the latest developments in GenAI.
@@ 3.3.3 Reliability evaluation
The reliability of the final item set was evaluated by focusing on internal consistency using Cronbach’s alpha. Cronbach’s alpha provides an estimate of the proportion of variance in the test scores attributable to the true score variance. A Cronbach’s alpha value of 0.7 or higher was deemed indicative of acceptable reliability in educational and psychological testing contexts, suggesting that the test items consistently assess the underlying construct of GenAI literacy (Miller, 1995). In addition, we utilised the coefficient omega, which is considered a more recent and potentially more accurate measure of internal consistency, particularly when items have varying loadings on the construct (Dunn et al., 2014). Specifically, we used omega total to evaluate the overall reliability of the GLAT in measuring GenAI literacy. This coefficient, like Cronbach’s alpha, ranges from 0 to 1 and utilises a similar threshold value (e.g., 0.7) to indicate acceptable reliability. Additionally, to further understand and confirm the precision and reliability of the GLAT, we evaluated the test information function (Reise and Waller, 2009). The test information function was analysed to ensure consistent measurement at proficiency levels most relevant to the GLAT’s intended application, particularly for learners and educators with low to average levels of GenAI literacy, considering GenAI technology is relatively new for them and in higher education (Annapureddy et al., 2024b; Jin et al., 2024).
3.4 RQ2: Incremental Validity
The incremental validity of the GLAT was assessed by analysing its predictive power concerning learners’ task performance during a task that involved interacting with a GenAI-powered conversational chatbot. This predictive capability was compared to that of a validated self-report instrument, the ChatGPT Literacy Scale (Lee and Park, 2024), to determine whether the GLAT provides additional predictive value beyond the self-report measure. The ChatGPT literacy scale serves as a suitable comparator, as it is specifically designed and validated for higher education students in the context of interactions with ChatGPT, a GenAI-powered conversational chatbot. Further details regarding the learning task, study procedure, and the analytical methods used are provided below.
@@ 3.4.1 Learning task
Learners engaged in a task that aims to enhance their ability to comprehend complex visual analytics, an ability that many are lacking (Maltese et al., 2015; Donohoe and Costello, 2020), through interacting with GenAI chatbots. The learning task involves learners analysing a set of visual analytics on students’ teamwork in healthcare simulations, composing a 100-word response and answering six evaluation questions to assess their ability to comprehend complex visual data. The visual analytics included three types of visualisations: a bar chart, a social network diagram, and a heatmap (Figure 1). Specifically, The bar chart illustrated students’ prioritisation strategies with positional data, simplifying the comparison of time spent on behaviours during the simulation [Anonymised]. The social network (sociogram) mapped interaction patterns via positional and audio data, highlighting communication frequencies and directions with the patient, doctor, and relative [Anonymised]. The advanced heatmap map combined students’ physical positions, verbal durations, and peak heart rate locations. Inspired by sports analytics (Goldsberry, 2012), it used heatmaps to show verbal communication frequency and spatial distribution, and identified areas of peak physiological arousal.
@@
Learners were instructed to first analyse the visualisations and write a 100-word response on how the two nurses managed the primary patient, Amy, while attending to other beds, focusing on task prioritisation, verbal communication, and stress levels. After this, they answered six multiple-choice questions (two per visualisation) designed to assess comprehension of the visual data, addressing the first two levels of Bloom’s taxonomy (knowledge and comprehension) (Bloom et al., 1984). For the knowledge questions, participants identified specific data points or patterns in visualisations, like determining which prioritisation behaviour two nurses spent the most time on from a bar chart. These questions assessed information retrieval skills. The comprehension questions required participants to interpret and derive insights, such as comparing spatial and verbal activities between two nurses using a ward map. These questions evaluated the ability to interpret insights and identify inconsistencies (see Table 4 for examples). Higher levels of Bloom’s taxonomy, like application, were not considered because the learners have limited contextual knowledge of visual analytics.
@@ 3.4.3 Study procedure and measures
As illustrated in Figure 3, learners begin by completing three literacy assessments: 1) the GLAT, 2) the ChatGPT Literacy Scale (Lee and Park, 2024), and 3) the mini-VLAT (Pandey and Ottley, 2023), which evaluates visualisation literacy pertinent to the task. Following a within-subject design, learners first perform the learning task independently, serving as the baseline condition. Subsequently, they repeat the task using different visual analytics (identical visualisations but with different data) while receiving support from the GenAI chatbot, constituting the AI-assisted condition. Consequently, five measures were captured and analysed: performance-based GenAI literacy using the GLAT (GLAT-literacy), self-reported GenAI literacy via the ChatGPT Literacy Scale (ChatGPT-literacy), visualisation literacy through the mini-VLAT (VLAT-literacy), baseline task performance without support (baseline score), and task performance with GenAI chatbot assistance (AI-assisted score).
@@ 4.1.4 Reliability
The reliability of the GLAT was assessed by examining internal consistency using Cronbach’s alpha and omega total. Cronbach’s alpha was calculated to be 0.80, indicating good reliability, as it surpasses the threshold of 0.7 commonly used in educational testing (Miller, 1995). In addition, the coefficient omega total was computed to further evaluate reliability, resulting in a value of 0.81. This suggests a high level of internal consistency and confirms the reliability of the GLAT in measuring GenAI literacy, particularly given the varying loadings on the construct (Dunn et al., 2014). Additionally, the precision and reliability of the GLAT were further examined using the test information function. As Figure 5 shows, the GLAT provides the most information at a proficiency level of $\theta=-0.8$, indicating that the test is highly reliable for individuals with below-average GenAI literacy. The maximum information value of 5.31 at this point suggests strong measurement precision for this target group. However, the information decreases for proficiency levels further from $\theta=-0.8$, especially for individuals with higher proficiency, where the test is less discriminative. The standard error (SE) was also lowest around $\theta=-0.8$, further supporting the test’s high precision for low to moderate GenAI literacy, while precision diminished for more extreme proficiency levels.
@@ 5 Discussion
Effective and valid measurement of GenAI literacy is essential in higher education as learners and educators increasingly encounter GenAI tools in their study, work, and daily lives (Yan et al., 2024a; Cukurova, 2024; Khosravi et al., 2023). This study developed and validated the GLAT in line with established standards for psychological and educational measurement (Thorndike et al., 1991; American Educational Research Association et al., 2014). Regarding RQ1, the GLAT demonstrated a 2PL model with strong structural validity, meeting the requirements of item discrimination and difficulty across a diverse sample. This indicates that the GLAT effectively differentiates individuals with varying GenAI literacy levels, which is crucial for accurately assessing competencies related to GenAI use (Annapureddy et al., 2024a; Yan et al., 2024a; Zhao et al., 2024a; Bozkurt, 2024a). The GLAT also showed good reliability, particularly in assessing students with low to moderate GenAI literacy. This aligns with its intended use, considering the current state of GenAI literacy, where the technology is relatively new and integrated training is limited in higher education curricula (Holmes et al., 2023; Jin et al., 2024). The GLAT is thus especially valuable for identifying individuals who may need additional education and support to effectively understand and use GenAI ethically. These findings highlight the GLAT’s utility in assessing foundational GenAI knowledge, particularly where students have limited prior exposure or training. However, as GenAI training becomes more integrated into higher education and students’ average GenAI literacy improves, the GLAT will require updates to remain relevant. This aligns with the need for an iterative design process for test instruments to adapt to new data and evolving use contexts (American Educational Research Association et al., 2014).
@@
In terms of RQ2, we examined the extent to which the GLAT predicts learners’ performance in tasks involving GenAI chatbots compared to self-reported instruments, using visualisation literacy and baseline performance as control variables. The predictive model showed that GenAI literacy, as measured by GLAT, was a significant predictor of learners’ performance in GenAI-supported tasks, whereas visualisation literacy served as a control to account for differences in comprehension of visual information. The significant positive relationship between GLAT and AI-assisted task performance underscores the value of objectively assessing GenAI literacy to predict real-world learner outcomes (Annapureddy et al., 2024a; Chiu, 2024a). In contrast, self-reported ChatGPT proficiency was not a significant predictor, highlighting the limitations of self-assessment, which may be prone to biases or inaccuracies (Lintner, 2024; Ng et al., 2021b). The control for visualisation literacy ensured that the observed effects were specific to GenAI literacy, thereby reinforcing the importance of targeted skill development in GenAI (Lee and Park, 2024; Lyu et al., 2024a). These findings suggest that enhancing GenAI literacy has a direct effect on learners’ ability to effectively engage with GenAI tools, independent of their visualisation skills. This insight is critical for educators aiming to design targeted interventions that bolster students’ competencies in using GenAI technologies effectively in diverse educational contexts.
@@ 5.1 Implications to Research and Practice
The study’s findings have profound implications for advancing research and practice in GenAI literacy within higher education. The development of the GLAT underscores the need for performance-based measures over traditional self-reported assessments, addressing the limitations and biases inherent in self-assessment tools (Ng et al., 2021b; Lintner, 2024). This shift to more reliable measures will enable educators and researchers to make informed decisions about integrating GenAI into educational settings. For educators, leveraging the GLAT offers a diagnostic tool to assess and enhance students’ GenAI literacy, highlighting individual needs for targeted interventions (Alzubi, 2024; Bozkurt, 2023). The assessment’s predictive validity in GenAI-supported tasks further underscores its practical utility, providing insights into students’ preparedness to navigate GenAI technologies in diverse educational contexts. By incorporating the GLAT into curriculum development, educators can better align teaching strategies with the specific competencies required for effective GenAI engagement, thereby preparing students for an AI-driven future. Furthermore, the study encourages researchers to adopt a comprehensive, iterative approach to developing and validating educational assessments, ensuring their continued relevance amidst the evolving GenAI landscape (Holmes et al., 2023; Jin et al., 2024). By focusing on multidimensional literacy frameworks that integrate foundational knowledge, practical skills, and ethical understanding, future research can enhance the robustness of educational instruments and cultivate a nuanced understanding of GenAI literacy in academia (Zhao et al., 2024b; Bozkurt, 2024a).

### v2

@@ 3 Methods
The GLAT was developed following the established test development procedures outlined in Psychological and Educational Measurement (Thorndike et al., 1991). The development process involved: 1) creating a blueprint of relevant GenAI concepts, 2) generating an initial set of test items based on this blueprint, and 3) evaluating face and content validity through expert reviews and pilot studies, respectively. An item analysis was conducted using CTT, which involved selecting items based on item difficulty and discrimination index. The structural validity and reliability of the GLAT (RQ1) were assessed using IRT. The external validity of the GLAT (RQ2) was evaluated by analysing its effectiveness in predicting learners’ performance on tasks involving interaction with a GenAI chatbot, compared to a self-reported GenAI literacy instrument (Lee and Park, 2024). Details were elaborated in the following sections.
@@ 3.1 Participants
Three samples of higher education students were involved in various validation processes for the GLAT, following established standards (Thorndike et al., 1991; American Educational Research Association et al., 2014). As shown in Figure 1, the first validation study involved assessing the content validity and selecting the GLAT item using CTT. Responses were gathered from 200 higher education students (101 females). The second validation study aimed to evaluate the structural validity and reliability of the GLAT using IRT (RQ1). This study analysed responses from 355 higher education students (184 females). The final validation study focused on assessing the external validity of the GLAT (RQ2) and included 83 higher education students (46 females). All participants were recruited through Prolific [^1] https://www.prolific.co/, a reputable online research recruitment platform. Participants in the first and second studies received £1.5 for their time, while those in the final study were compensated £8 due to its increased complexity, as elaborated in Section 3.4. All studies were conducted using Qualtrics, with the item and option orders in the GLAT randomised (American Educational Research Association et al., 2014). Ethics approval was obtained from [Anonymised] University (Project ID: Anonymised), and informed consent was obtained from all participants.
Figure 1: The participant sample size and focus of each validation study.
@@ 3.2.1 Blueprint construction
The blueprint of the GLAT was developed based on the four dimensions of AI literacy proposed by Ng et al. (2021b), including 1) Know & Understand, 2) Use & Apply, 3) Evaluate & Create, and 4) Ethics, and focusing on the specific context of GenAI. To identify a set of relevant GenAI concepts in each dimension, we cover a wide range of resources, including academic publications in prestige journals (e.g., Nature and Science), reports and articles published by reputable organisations (e.g., UNESCO, MIT News, and Standford HAI), and education information released by leading GenAI technology companies (e.g., OpenAI, Google, Meta, and NVIDIA). The decision to include diverse sources, beyond traditional academic publications, was driven by the rapidly evolving landscape of GenAI. This encompasses foundational models (e.g., large language models and diffusion models) as well as supporting infrastructure and techniques (e.g., embedding databases and retrieval methods). Our process for extracting relevant concepts involved three steps. Initially, two researchers independently reviewed the source documents (n=19; links are available in the repository) and recorded pertinent GenAI concepts. Subsequently, they collaborated to consolidate similar concepts, resulting in a refined set of 25 concepts (see Table 1). A validation panel of three GenAI researchers then reviewed these concepts to ensure they provided reasonable coverage of the latest developments in GenAI.
@@ 3.3.3 Reliability evaluation
The reliability of the final item set was evaluated by focusing on internal consistency using Cronbach’s alpha. Cronbach’s alpha provides an estimate of the proportion of variance in the test scores attributable to the true score variance. A Cronbach’s alpha value of 0.7 or higher was deemed indicative of acceptable reliability in educational and psychological testing contexts, suggesting that the test items consistently assess the underlying construct of GenAI literacy (Miller, 1995). In addition, we utilised the coefficient omega, which is considered a more recent and potentially more accurate measure of internal consistency, particularly when items have varying loadings on the construct (Dunn et al., 2014). Specifically, we used omega total to evaluate the overall reliability of the GLAT in measuring GenAI literacy. This coefficient, like Cronbach’s alpha, ranges from 0 to 1 and utilises a similar threshold value (e.g., 0.7) to indicate acceptable reliability. Additionally, to further understand and confirm the precision and reliability of the GLAT, we evaluated the test information function (Reise and Waller, 2009). The test information function was analysed to ensure consistent measurement at proficiency levels most relevant to the GLAT’s intended application, particularly for learners and educators with low to average levels of GenAI literacy, considering GenAI technology is relatively new for them and in higher education (Annapureddy et al., 2024; Jin et al., 2024).
3.4 RQ2: External Validity
The external validity of the GLAT was assessed by analysing its predictive power concerning learners’ task performance during a task that involved interacting with a GenAI-powered conversational chatbot. The choice of a task involving interaction with a GenAI-powered chatbot is particularly appropriate for assessing external validity, as it closely mirrors real-world applications and challenges students might encounter in educational settings (McGrath et al., 2024), thereby providing a practical context for evaluating the external validity of GenAI literacy in predicting learner performance. This predictive capability was compared to that of a validated self-report instrument, the ChatGPT Literacy Scale (Lee and Park, 2024), to determine whether the GLAT provides additional predictive value beyond the self-report measure. The ChatGPT literacy scale serves as a suitable comparator, as it is specifically designed and validated for higher education students in the context of interactions with ChatGPT, a GenAI-powered conversational chatbot. Further details regarding the learning task, study procedure, and the analytical methods used are provided below.
@@ 3.4.1 Learning task
Learners engaged in a task that aimed to enhance their ability to comprehend complex visual analytics, an ability that many are lacking (Maltese et al., 2015; Donohoe and Costello, 2020), through interacting with GenAI chatbots. The learning task involved learners analysing a set of visual analytics on students’ teamwork in healthcare simulations, composing a 100-word response and answering six evaluation questions to assess their ability to comprehend complex visual data. The visual analytics included three types of visualisations: a bar chart, a social network diagram, and a heatmap (Figure 2). Specifically, the bar chart illustrated students’ prioritisation strategies with positional data, simplifying the comparison of time spent on behaviours during the simulation [Anonymised]. The social network (sociogram) mapped interaction patterns via positional and audio data, highlighting communication frequencies and directions with the patient, doctor, and relative [Anonymised]. The advanced heatmap map combined students’ physical positions, verbal durations, and peak heart rate locations. Inspired by sports analytics (Goldsberry, 2012), it used heatmaps to show verbal communication frequency and spatial distribution, and identified areas of peak physiological arousal.
@@
A total of 83 higher education students (46 females) with medical, healthcare, and nursing backgrounds were involved in this validation to control for their familiarity with the healthcare simulation context. Learners were instructed to first analyse the visualisations and write a 100-word response on how the two nurses managed the primary patient, Amy, while attending to other beds, focusing on task prioritisation, verbal communication, and stress levels. After this, they answered six multiple-choice questions (two per visualisation) designed to assess comprehension of the visual data, addressing the first two levels of Bloom’s taxonomy (knowledge and comprehension) (Bloom et al., 1984). For the knowledge questions, participants identified specific data points or patterns in visualisations, like determining which prioritisation behaviour two nurses spent the most time on from a bar chart. These questions assessed information retrieval skills. The comprehension questions required participants to interpret and derive insights, such as comparing spatial and verbal activities between two nurses using a ward map. These questions evaluated the ability to interpret insights and identify inconsistencies (see Table 4 for examples). Higher levels of Bloom’s taxonomy, like application, were not considered because the learners have limited contextual knowledge of visual analytics.
@@ 3.4.3 Study procedure and measures
As illustrated in Figure 4, learners begin by completing three literacy assessments: 1) the GLAT, 2) the ChatGPT Literacy Scale (Lee and Park, 2024), and 3) the mini-VLAT (Pandey and Ottley, 2023), which evaluates domain knowledge, specifically, learners’ visualisation literacy pertinent to the task. Following a within-subject design, learners first perform the learning task independently, serving as the baseline condition. Subsequently, they repeat the task using different visual analytics (identical visualisations but with different data) while receiving support from the GenAI chatbot, constituting the AI-assisted condition. Consequently, five measures were captured and analysed: performance-based GenAI literacy using the GLAT (GLAT-literacy), self-reported GenAI literacy via the ChatGPT Literacy Scale (ChatGPT-literacy), visualisation literacy through the mini-VLAT (VLAT-literacy), baseline task performance without support (baseline score), and task performance with GenAI chatbot assistance (AI-assisted score).
@@ 4.1.4 Reliability
The reliability of the GLAT was assessed by examining internal consistency using Cronbach’s alpha and omega total. Cronbach’s alpha was calculated to be 0.80, indicating good reliability, as it surpassed the threshold of 0.7 commonly used in educational testing (Miller, 1995). In addition, the coefficient omega total was computed to further evaluate reliability, resulting in a value of 0.81. This suggested a high level of internal consistency and confirmed the reliability of the GLAT in measuring GenAI literacy, particularly given the varying loadings on the construct (Dunn et al., 2014). Additionally, the precision and reliability of the GLAT were further examined using the test information function. As Figure 6 shows, the GLAT provided the most information at a proficiency level of $\theta=-0.8$, indicating that the test is highly reliable for individuals with below-average GenAI literacy. The maximum information value of 5.31 at this point suggests strong measurement precision for this target group. However, the information decreases for proficiency levels further from $\theta=-0.8$, especially for individuals with higher proficiency, where the test is less discriminative. The standard error (SE) was also lowest around $\theta=-0.8$, further supporting the test’s high precision for low to moderate GenAI literacy, while precision diminished for more extreme proficiency levels.
@@ 5 Discussion
Effective and valid measurement of GenAI literacy is essential in higher education as learners and educators increasingly encounter GenAI tools in their study, work, and daily lives (Yan et al., 2024a; Cukurova, 2024; Khosravi et al., 2023). This study developed and validated the GLAT in line with established standards for psychological and educational measurement (Thorndike et al., 1991; American Educational Research Association et al., 2014). Regarding RQ1, the GLAT demonstrated a 2PL model with strong structural validity, meeting the requirements of item discrimination and difficulty across a diverse sample. This indicates that the GLAT effectively differentiates individuals with varying GenAI literacy levels, which is crucial for accurately assessing competencies related to GenAI use (Annapureddy et al., 2024; Yan et al., 2024a; Zhao et al., 2024; Bozkurt, 2024a). The GLAT also showed good reliability, particularly in assessing students with low to moderate GenAI literacy. This aligns with its intended use, considering the current state of GenAI literacy, where the technology is relatively new and integrated training is limited in higher education curricula (Holmes et al., 2023; Jin et al., 2024). The GLAT is thus especially valuable for identifying individuals who may need additional education and support to effectively understand and use GenAI ethically. These findings highlight the GLAT’s utility in assessing foundational GenAI knowledge, particularly where students have limited prior exposure or training. However, as GenAI training becomes more integrated into higher education and students’ average GenAI literacy improves, the GLAT will require updates to remain relevant. This aligns with the need for an iterative design process for test instruments to adapt to new data and evolving use contexts (American Educational Research Association et al., 2014).
@@
In terms of RQ2 and external validity, we examined the extent to which the GLAT predicts learners’ performance in tasks involving GenAI chatbots compared to self-reported instruments, using visualisation literacy and baseline performance as control variables. The predictive model showed that GenAI literacy, as measured by GLAT, was a significant predictor of learners’ performance in GenAI-supported tasks, whereas domain knowledge (e.g., visualisation literacy) served as a control to account for differences in learners’ comprehension of visual information. The significant positive relationship between GLAT and AI-assisted task performance underscores the value of reliably assessing GenAI literacy to predict real-world learner outcomes (Annapureddy et al., 2024; Chiu, 2024). In contrast, self-reported ChatGPT proficiency was not a significant predictor, highlighting the limitations of self-assessment, which may be prone to biases or inaccuracies (Lintner, 2024; Ng et al., 2021b). The control for domain knowledge (e.g., visualisation literacy) ensured that the observed effects were specific to GenAI literacy, thereby reinforcing the importance of targeted skill development in GenAI (Lee and Park, 2024; Lyu et al., 2024). These findings suggest that enhancing GenAI literacy has a direct effect on learners’ ability to effectively engage with GenAI tools, independent of their domain knowledge. This insight is critical for educators aiming to design targeted interventions that bolster students’ competencies in using GenAI technologies effectively in diverse educational contexts.
@@ 5.1 Implications to Research and Practice
The study’s findings have profound implications for advancing research and practice in GenAI literacy within higher education. The development of the GLAT underscores the need for performance-based measures over traditional self-reported assessments, addressing the limitations and biases inherent in self-assessment tools (Ng et al., 2021b; Lintner, 2024). This shift to more reliable measures will enable educators and researchers to make informed decisions about integrating GenAI into educational settings. For educators, leveraging the GLAT offers a diagnostic tool to assess and enhance students’ GenAI literacy, highlighting individual needs for targeted interventions (Alzubi, 2024; Bozkurt, 2023). The assessment’s external validity in GenAI-supported learning tasks further underscores its practical utility, providing insights into students’ preparedness to navigate GenAI technologies in diverse educational contexts. By incorporating the GLAT into curriculum development, educators can better align teaching strategies with the specific competencies required for effective GenAI engagement, thereby preparing students for an AI-driven future. Furthermore, the study encourages researchers to adopt a comprehensive, iterative approach to developing and validating educational assessments, ensuring their continued relevance amidst the evolving GenAI landscape (Holmes et al., 2023; Jin et al., 2024). By focusing on multidimensional literacy frameworks that integrate foundational knowledge, practical skills, and ethical understanding, future research can enhance the robustness of educational instruments and cultivate a nuanced understanding of GenAI literacy in academia (Zhao et al., 2024; Bozkurt, 2024a).